OpenStudy (anonymous):

write y=3/7x-2 in standard form w/ integer coefficients.

OpenStudy (lalaly):

multiply both sides by 7 7y=3x-14 3x-7y-14=0

OpenStudy (anonymous):

Standard form is ax+by=c. Where all values are positive integers. So subtract the x term to the other side. -3/7x+y=-2 Now, eliminate fractions by multiplying by the LCD (7) -3x+7y=-14 Sometimes the answer will have a positive first term instead, so you can multiply everything by -1 if you'd like to. 3x-7y=14